A) On January 1, 2018, Cullumber Corp. had 480,000 shares of common stock outstanding. During 2018, it had the following transactions that affected the Common Stock account.

February 1 Issued 117,000 shares

March 1 Issued a 10% stock dividend

May 1 Acquired 101,000 shares of treasury stock

June 1 Issued a 3-for-1 stock split

October 1 Reissued 62,000 shares of treasury stock

Determine the weighted-average number of shares outstanding as of December 31, 2018. The weighted-average number of shares outstanding

B) Assume that Cullumber Corp. earned net income of $3,500,000 during 2018. In addition, it had 96,000 shares of 9%, $100 par nonconvertible, noncumulative preferred stock outstanding for the entire year. Because of liquidity considerations, however, the company did not declare and pay a preferred dividend in 2018. Compute earnings per share for 2018, using the weighted-average number of shares determined in part (a). (Round answer to 2 decimal places, e.g. $2.55.)

Earnings Per Share $

C) Assume the same facts as in part (b), except that the preferred stock was cumulative. Compute earnings per share for 2018. (Round answer to 2 decimal places, e.g. $2.55.)

Earnings Per Share $

D) Assume the same facts as in part (b), except that net income included a loss from discontinued operations of $452,000 (net of tax). Compute earnings per share for 2018.
Answer #1

A) Calculation of weighted average number of share outstanding -

Formula for calculating the same is = Number of share outstanding X (number of months share outstanding)/(total number of months)

480,000 X 12/12 = 480,000

117,000 x 11/12 = 107,250

59,700 x 10/12 = 49,750 (Refer WN 1)

555,700 x 8/12 = 370,466 (Refer WN 2)

1,667,100 x 7/12 = 972,475

62,000 x 3/12 = 15500

Therefore, total number of weighted average shares outstanding is = (480,000+107,250+49,750+370,466+972,475+15,500) = 1,995,441

W.N 1 Stock Dividend = (480000+117000) X 10% = 59,700

W.N. 2 Acquired 101,000 treasurry stock. Therefore stock left in the market = (480,000+117000+59700-101000) = 555,700

W.N. 3 Total number of stock in the market after acquisition = 555700

After stock split, total number of shares left = 555700 x 3 = 1,667,100

  

B) Net Income for 2018 = 3,500,000

Total W A Number of shares outstanding = 1,995,441

Earning per share (EPS) = 3,500,000/1,995,441 = 1.75

(Earning per share, when preferred dividend is not paid = Net Income / Total WA Number of shares outstanding).

C) If the preferred stock is cumulative, then net income will be = 3,500,000 - (96,000 X 100 X 9%) = 2,636,000.

Now the EPS will be 2,636,000/1,995,441 = 1.32.

D) Net income will be = (3,500,000 - 4,52,000) = 3,048,000. (After adjusting the loss from discontinued operations)

EPS will be 3,048,000/1,995,441 = 1.53